# Searching for local associations while controlling the false discovery rate

### v1

@@ 1.2 Main Contributions and Outline
This approach introduces a new way to efficiently partition the information in a dataset, allowing both learning and testing (Tian and Taylor, 2018; Rasines and Young, 2023; Leiner et al., 2023). Thus, our method can be seen as a “data-thinning” approach (Dharamshi et al., 2024), specifically designed for knockoff inference and taking advantage of the unique capacity of knockoffs to partially “mask” data (Spector and Fithian, 2022).
@@ 1.3 Related Work
This paper focuses on identifying subgroups of a heterogeneous population within which there are relatively homogeneous conditional associations. There are many algorithms and models that can discover subgroups of individuals exhibiting coherent behaviors (Novak et al., 2009), detect interactions (Bien et al., 2013; Wen and Stephens, 2014; Tibshirani and Friedman, 2020), or estimate heterogeneous effects in regression and causal inference (Athey and Imbens, 2016; Moore et al., 2019; Künzel et al., 2019; Hahn et al., 2020; Cui et al., 2023); see also Caron et al. (2022) for a review. While such methods are usually not designed to obtain statistical inferences with finite-sample guarantees, they provide powerful tools that can be leveraged in the “learning” phase of our method.
This work is also inspired by the literature on explainability in machine learning; e.g., see Christoph (2020) for a review. A central challenge in this field is enhancing the interpretability of opaque predictive models by identifying “important” variables. There are many definitions of variable importance, from descriptive measures capturing a model’s input-output relationship to inferential measures offering insights into the underlying population. Given the nonlinear nature of many machine learning models, assessments of variable importance are often context-dependent. Despite limitations in current approaches (Ghassemi et al., 2021; Mase et al., 2024), this research highlights the value of “local” importance measures in capturing data heterogeneity.
@@ 2.1 Review of Global Conditional Hypotheses
Consider for example a simple mixture population with $p=2$ and $m=1$, with $Z_{1}=1$ for half of the individuals and $Z_{1}=0$ for the others, indicating two separate environments. Suppose $X_{1}$ and $X_{2}$ are independent standard normal and $Y\sim\mathcal{N}(Z_{1}X_{2}\beta_{2},1)$, for some coefficient $\beta_{2}\in\mathbb{R}$. The hypotheses in (1) are well defined, with $\mathcal{H}_{1}$ true and $\mathcal{H}_{2}$ true if and only if $\beta_{2}=0$. Thus, testing (1) is already useful for distinguishing the importance of $X_{2}$ in predicting $Y$ from the irrelevance of $X_{1}$. However, one limitation of this approach, which we address in this paper, is that it does not inform about how the importance of $X_{2}$ varies with changes in the value of $Z_{1}$. In other words, testing (1) cannot tells us that $X_{2}$ is only locally important among the subgroup with $Z_{1}=1$.
@@ 2.2 A Model-X Framework for Heterogeneous Data
where each $p^{i}_{Y\mid X,Z}$ denotes the (unknown) conditional distribution of the outcome $Y^{i}$ for the $i$-th individual, given that individual’s variables and covariates. This distribution may vary across individuals, who are assumed to be independent. To simplify notation, we assume that $p_{X,Z}$ remains constant, though this assumption could be relaxed. The concept of an individual-specific distribution $p^{i}_{Y\mid X,Z}$ is abstract and provides limited opportunities for learning, as each individual essentially forms their own category. We use this framework to capture unknown heterogeneity, but in practice, learning will be conducted at the group level by aggregating information from “similar” individuals, as it will soon become clear.
@@ 2.4 Modeling Assumptions and Knockoffs
In general, Equation (3) can be factored as $P(\mathbf{X},\mathbf{Y},\mathbf{Z})=\prod_{i=1}^{n}p_{Z}(Z^{i})\cdot p_{X\mid Z}(Z^{i}\mid X^{i})\cdot p^{i}_{Y\mid X,Z}(Y^{i}\mid X^{i},Z^{i})$. Following a model-X approach (Candès et al., 2018), we treat $\mathbf{Z}$ as fixed and $P_{X\mid Z}$ as known, while allowing $p^{i}_{Y\mid X,Z}$ to be arbitrary and unknown for all $i\in[n]$.
@@
Constructing knockoffs that satisfy (7) not only requires knowledge of $P_{X\mid Z}$ but it can also be computationally expensive in general. However, the problem is well-studied (Candès et al., 2018; Gimenez et al., 2019; Bates et al., 2020), with practical and effective solutions available for GWAS data (Sesia et al., 2019, 2020, 2021). Therefore, in this paper we assume the availability of suitable knockoffs for the data at hand, using this as a starting point for developing our methodology to test local conditional independence hypotheses.
@@ 2.5 A Knockoff Filter for Fixed Local Hypotheses
To enable provable FDR control, these scores must be computed in such a way that swapping a variable $X_{j}$ with its knockoff $\tilde{X}_{j}$ within any subgroup $\ell\in[L_{j}]$ would have the sole effect of swapping the corresponding scores and in that subgroup, namely $T_{j,\ell}$ and $\tilde{T}_{j,\ell}$. We now formalize this principle within a general framework, which not only encompasses all implementations of our method applied in this paper but also accommodates a range of potential future extensions. This framework is based on a data augmentation idea.
Let $\mathbf{V}\in\{0,1\}^{n\times p}$ be i.i.d. Bernoulli random variables, representing a random noise matrix generated by the analyst independent of everything else. Then, let $[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V})}\in\mathbb{R}^{n\times 2p}$ denote the concatenation of $\mathbf{X}$ and $\tilde{\mathbf{X}}$, with the $i$-th observation of $X_{j}$ swapped with its knockoff if and only if $V_{ij}=1$. Combining $[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V})}$ with $\mathbf{Y}$ and $\mathbf{Z}$ gives us a cloaked data set $\tilde{\mathcal{D}}(\mathbf{V})=(\smash{[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V})},\mathbf{Y},\mathbf{Z}})$, where the identities of the original variables and knockoffs are hidden. Although $\tilde{\mathcal{D}}(\mathbf{V})$ (deliberately) contains additional noise compared to $\mathcal{D}$, it still carries valuable information about the relation between $X$ and $Y$ because each column of the matrix $\smash{[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V})}}$ contains approximately $n/2$ real observations.
@@
One way to achieve (9) is to compute each pair $(T_{j,\ell},\tilde{T}_{j,\ell})$ using an appropriate measure of the importance of $(X_{j},\tilde{X}_{j})$ within a local model $\hat{\psi}^{(j,\ell)}$ predicting $Y$ given $(X,\tilde{X},Z)$. This model can be trained using the cloaked dataset $\tilde{\mathcal{D}}(\mathbf{V})$ and a smaller local dataset $\mathcal{D}^{(j,\ell)}=(\mathbf{X}_{j},\tilde{\mathbf{X}}_{j},\mathbf{Y},\mathbf{Z})^{(j,\ell)}$, including only observations $(X_{j},\tilde{X}_{j},Y,Z)$ for individuals $i\in[n]$ with $\nu_{j}(Z^{i})=\ell$. Figure A2 in Appendix A2 provides a schematic visualization of this workflow. Then, to guarantee (9), it suffices to compute $(T_{j,\ell},\tilde{T}_{j,\ell})$ in such a way that swapping $X_{j}$ and $\tilde{X}_{j}$ in $\mathcal{D}^{(j,\ell)}$ results in $T_{j,\ell}$ being swapped with $\tilde{T}_{j,\ell}$; see Proposition A3 in Appendix A3. With a slight abuse of notation, we write this local symmetry property as:
@@ 2.6 A Knockoff Filter for Data-Driven Local Hypotheses
This two-step framework for adaptively selecting and testing local hypotheses is outlined by Algorithm 2 and visualized schematically in Figure A2 in Appendix A2. We call this approach Adaptive Local Knockoff Filter (aLKF).
@@ 4 Apply Algorithm 1 using the partition function $\hat{\nu}$;
For example, we can fit the lasso to predict $\mathbf{Y}$ using $\smash{[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V})}}$ and $\smash{\mathbf{Z}_{1}}$, including all possible pairwise interactions between $\mathbf{Z}_{1}$ and $\smash{[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V})}}$. For each $j\in[p]$, the partition function $\smash{\hat{\nu}_{j}}$ then divides the samples into two subgroups, corresponding to $\nu_{j}(0)=1$ and $\nu_{j}(1)=1$, if the model selects at least one interaction term for either the $j$-th or $(j+p)$-th variables in $\smash{[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V})}}$. Otherwise, $\smash{\hat{\nu}_{j}(z)}=1$ for all $z$ values. In general, when dealing with $m$ covariates, this approach is applied using a hyper-parameter $G_{\max}\in\mathbb{N}$ controlling the maximum number of interactions that may be selected per variable (e.g., $G_{\max}=2$).
As detailed in Appendix A2.2, this implementation can be extended to handle multiple binary covariates as well as non-binary covariates. Additionally, Appendix A2.3 describes a specialized implementation with computational shortcuts designed for GWAS-like settings, characterized by a very large number $p$ of variables and a modest number $m$ of covariates.
@@ 3 Empirical Demonstrations
We demonstrate the performance of aLKF in two simulated settings and one data analysis.
@@ 3.1 Simulation: Finding Highly Localized Important Variables
We generate synthetic data with $p=20$ binary variables and $m=80$ covariates, where the first 20 covariates are binary and the remaining 60 are continuous in $[0,1]$. Details on the joint distribution of these variables and the construction of knockoffs are in Appendix A4.
@@
Given a sample of $n$ independent observations, our goal is to identify which variables are non-null within which subgroups, aiming for both power (many rejections) and precision (informative subgroups). We apply our method (aLKF) at the nominal FDR level $\alpha=0.1$, using the lasso-based approach for importance scores (Appendix A2.1) and learning data-driven partitions via a lasso model with interactions (Appendix A2.2), allowing at most two interactions per variable ($G_{\max}=2$). We compare aLKF against three benchmarks: (1) Global-KF, the knockoff filter of Candès et al. (2018), which controls the FDR but tests only population-wide hypotheses; (2) Split-LKF, a less powerful version of aLKF that splits the data, using half to learn the partition function and the other half to test local hypotheses, sacrificing power due to reduced sample size but maintaining FDR control; (3) Naive-LKF, an invalid version of aLKF that uses all data twice—to learn the partition and to test local hypotheses—without cloaking, failing to control the FDR due to selection bias.
@@ 3.2 Simulation: Searching for Gene-Environment Interactions
The following experiments simulate a search for gene-environment interactions in GWAS data, using the real chromosome-21 genotypes of unrelated white non-British subjects in the UK Biobank. We use the knockoffs defined by Sesia et al. (2021). There are $14,733$ individuals in total and $p=8,832$ variants. We randomly sample 30% of the individuals ($n=4,420$) to imitate a more typical high-dimensional GWAS setting.
We simulate two independent binary random covariates $Z_{1}$ and $Z_{2}$, independently across individuals from a Bernoulli($0.5$) distribution. The results of additional experiments using only one covariate can be found in Appendix A6. A continuous outcome $Y$ is simulated based on a linear model: $Y^{i}=\beta_{0}X^{i}+(\beta_{1}X^{i})*\mathbf{1}(Z_{1}^{i}=0)+\epsilon^{i}$, with $\beta_{0},\beta_{1}\in\mathbb{R}^{p}$, where $\epsilon^{i}\sim N(0,1)$ independently for each $i\in[n]$. Thus, $Z_{1}$ is used to create local effects, among the subgroup with $Z_{1}=0$, while $Z_{2}$ is unimportant. We randomly choose 1% of the genetic variants to be non-null, in the sense that they have a non-zero coefficient either globally (in $\beta_{0}$) or locally (in $\beta_{1}$). Additional results obtained using different proportions of non-null variants (0.05% and 2%) can be found in Appendix A6. Each non-null variant is randomly assigned as either a global or a local effect according to the desired proportions.
@@
To maintain sufficient power in the presence of strong collinearity among nearby genetic variants on the same chromosome (linkage disequilibrium), we adopt the approach of Sesia et al. (2020, 2021). The genome is partitioned into contiguous blocks of variants in strong linkage disequilibrium, with conditional hypotheses tested at the block level. These blocks serve as the units of inference, replacing individual SNPs. In this simulation, the median width of the blocks is 3 kilobases (kb); see Appendices A2.3.4 and A5 for further details.
Our goal is to identify which variables are non-null within which subgroups, seeking both high power and precision. We apply aLKF as described in Section 3.1, allowing the data-driven partition function to select at most one interaction per variable ($G_{\max}=1$). In addition to the three benchmarks from Section 3.1Global-KF, Split-LKF, and Naive-LKF—we introduce a fourth benchmark, Fixed-LKF, a non-adaptive version of LKF.

### v2

@@ 1.2 Main Contributions and Outline
This method efficiently partitions the information in a dataset, allowing both learning and testing (Tian and Taylor, 2018; Leiner et al., 2023). Thus, it can be seen as a “data-thinning” approach (Dharamshi et al., 2024), specifically designed for knockoff inference and taking advantage of the unique capacity of knockoffs to partially “mask” data (Spector and Fithian, 2022).
@@ 1.3 Related Work
This paper focuses on identifying subgroups of a heterogeneous population within which there are relatively homogeneous conditional associations. There are many algorithms and models that can discover subgroups exhibiting coherent behaviors (Novak et al., 2009; Chernozhukov et al., 2018), detect interactions (Bien et al., 2013; Wen and Stephens, 2014; Tibshirani and Friedman, 2020), or estimate heterogeneous effects in regression and causal inference (Athey and Imbens, 2016); see also Caron et al. (2022) for a review. While such methods are usually not designed to obtain statistical inferences with finite-sample guarantees, they can be leveraged in the “learning” phase of our method.
This work is also inspired by the literature on explainability in machine learning (Molnar, 2020), where a central challenge in this field is enhancing the interpretability of opaque predictive models by identifying “important” variables. There are many definitions of variable importance, from descriptive measures capturing a model’s input-output relationship to inferential measures offering insights into the underlying population. Given the nonlinear nature of many machine learning models, assessments of variable importance are often context-dependent. Despite limitations in current approaches (Ghassemi et al., 2021; Mase et al., 2024), this research highlights the value of “local” importance measures in capturing data heterogeneity.
@@ 2.1 Review of Global Conditional Hypotheses
Consider for example a simple mixture population with $p=2$ and $m=1$, with $Z_{1}=1$ for half of the individuals and $Z_{1}=0$ for the others, indicating two separate environments. Suppose $X_{1}$ and $X_{2}$ are independent standard normal and $Y\sim\mathcal{N}(Z_{1}X_{2}\beta_{2},1)$, for some coefficient $\beta_{2}\in\mathbb{R}$. The hypotheses in (1) are well defined, with $\mathcal{H}_{1}$ true and $\mathcal{H}_{2}$ true if and only if $\beta_{2}=0$. Thus, testing (1) is already useful for distinguishing the importance of $X_{2}$ in predicting $Y$ from the irrelevance of $X_{1}$. However, one limitation of this approach, which we address in this paper, is that it does not inform about how the importance of $X_{2}$ varies with changes in the value of $Z_{1}$. In other words, testing (1) cannot tell us that $X_{2}$ is only locally important among the subgroup with $Z_{1}=1$.
@@ 2.2 A Model-X Framework for Heterogeneous Data
where each $p^{i}_{Y\mid X,Z}$ denotes the (unknown) conditional distribution of the outcome $Y^{i}$ for the $i$-th individual, given that individual’s variables and covariates. This distribution may vary across individuals, who are assumed to be independent. To simplify notation, we assume that $p_{X,Z}$ remains constant, though this assumption could be relaxed. The concept of an individual-specific distribution $p^{i}_{Y\mid X,Z}$ is abstract and provides limited opportunities for learning, as each individual essentially forms their own category. We use this framework to capture unknown heterogeneity, but in practice, learning will be conducted at the subgroup level by aggregating information from “similar” individuals, as it will soon become clear.
@@ 2.4 Modeling Assumptions and Knockoffs
In general, Equation (3) can be factored as $P(\mathbf{X},\mathbf{Y},\mathbf{Z})=\prod_{i=1}^{n}p_{Z}(Z^{i})\cdot p_{X\mid Z}(X^{i}\mid Z^{i})\cdot p^{i}_{Y\mid X,Z}(Y^{i}\mid X^{i},Z^{i})$. Following a model-X approach (Candès et al., 2018), we treat $\mathbf{Z}$ as fixed and $P_{X\mid Z}$ as known, while allowing $p^{i}_{Y\mid X,Z}$ to be arbitrary and unknown for all $i\in[n]$.
@@
Constructing knockoffs that satisfy (7) not only requires knowledge of $P_{X\mid Z}$ but it can also be computationally expensive in general. However, the problem is well-studied (Candès et al., 2018; Gimenez et al., 2019; Bates et al., 2020), with practical and effective solutions available for GWAS data (Chu et al., 2024; Sesia et al., 2019, 2020, 2021). Therefore, in this paper we assume the availability of suitable knockoffs for the data at hand, using this as a starting point for developing our methodology to test local conditional independence hypotheses. In our GWAS analyses specifically, we follow the approach of Sesia et al. (2021) to generate valid knockoffs, leveraging hidden Markov models.
@@ 2.5 A Knockoff Filter for Fixed Local Hypotheses
To enable provable FDR control, these scores must be computed in such a way that swapping a variable $X_{j}$ with its knockoff $\tilde{X}_{j}$ within any subgroup $\ell\in[L_{j}]$ would have the sole effect of swapping the corresponding scores in that subgroup, namely $T_{j,\ell}$ and $\tilde{T}_{j,\ell}$. We now formalize this principle within a general framework, which not only encompasses all implementations of our method applied in this paper but also accommodates a range of potential future extensions. This framework is based on a data augmentation idea.
Let $\mathbf{V}\in\{0,1\}^{n\times p}$ be i.i.d. Bernoulli(0.5) random variables, representing a random noise matrix generated by the analyst independent of everything else. Then, let $[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V})}\in\mathbb{R}^{n\times 2p}$ denote the concatenation of $\mathbf{X}$ and $\tilde{\mathbf{X}}$, with the $i$-th observation of $X_{j}$ swapped with its knockoff if and only if $V_{ij}=1$. Combining $[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V})}$ with $\mathbf{Y}$ and $\mathbf{Z}$ gives us a cloaked data set $\tilde{\mathcal{D}}(\mathbf{V})=(\smash{[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V})},\mathbf{Y},\mathbf{Z}})$, where the identities of the original variables and knockoffs are hidden. Although $\tilde{\mathcal{D}}(\mathbf{V})$ (deliberately) contains additional noise compared to $\mathcal{D}$, it still carries valuable information about the relation between $X$ and $Y$ because each column of the matrix $\smash{[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V})}}$ contains approximately $n/2$ real observations.
@@
One way to achieve (9) is to compute each pair $(T_{j,\ell},\tilde{T}_{j,\ell})$ using an appropriate measure of the importance of $(X_{j},\tilde{X}_{j})$ within a local model $\hat{\psi}^{(j,\ell)}$ predicting $Y$ given $(X,\tilde{X},Z)$. This model can be trained using the cloaked dataset $\tilde{\mathcal{D}}(\mathbf{V})$ and a smaller local dataset $\mathcal{D}^{(j,\ell)}=(\mathbf{X}_{j},\tilde{\mathbf{X}}_{j},\mathbf{Y},\mathbf{Z})^{(j,\ell)}$, including only observations $(X_{j},\tilde{X}_{j},Y,Z)$ for individuals $i\in[n]$ with $\nu_{j}(Z^{i})=\ell$. Figure 1 provides a schematic visualization of this workflow. Then, to guarantee (9), it suffices to compute $(T_{j,\ell},\tilde{T}_{j,\ell})$ in such a way that swapping $X_{j}$ and $\tilde{X}_{j}$ in $\mathcal{D}^{(j,\ell)}$ results in $T_{j,\ell}$ being swapped with $\tilde{T}_{j,\ell}$; see Proposition A3 in Appendix A3. With a slight abuse of notation, we write this local symmetry property as:
@@ 2.6 A Knockoff Filter for Data-Driven Local Hypotheses
This two-step framework for adaptively selecting and testing local hypotheses is outlined by Algorithm 2 and visualized schematically in Figure 1 below. We call this approach Adaptive Local Knockoff Filter (aLKF).
Figure 1: Schematic of the Adaptive Local Knockoff Filter. The relevant variables and associated statistics for different local environments are shown in different colors. The original variables are shown in darker shades and the knockoffs in lighter shades.
@@ 4 Apply Algorithm 1 using the partition function $\hat{\nu}$;
For example, we can fit the lasso to predict $\mathbf{Y}$ using $\smash{[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V})}}$ and $\smash{\mathbf{Z}_{1}}$, including all possible pairwise interactions between $\mathbf{Z}_{1}$ and $\smash{[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V})}}$. For each $j\in[p]$, the partition function $\smash{\hat{\nu}_{j}}$ then divides the samples into two subgroups, corresponding to $\nu_{j}(0)=1$ and $\nu_{j}(1)=1$, if the model selects at least one interaction term for either the $j$-th or $(j+p)$-th variables in $\smash{[\mathbf{X},\tilde{\mathbf{X}}]_{\mathrm{swap}(\mathbf{V})}}$. Otherwise, $\smash{\hat{\nu}_{j}(z)}=1$ for all $z$ values. In general, when dealing with $m$ covariates, this approach is applied using a hyper-parameter $G_{\max}\in\mathbb{N}$ controlling the maximum number of interactions that may be selected per variable. The choice of $G_{\max}$ does not affect the theoretical validity of our method, but it mediates a trade-off: larger values may reveal more specific (and thus potentially more homogeneous) subgroups, but can also reduce power to reject the associated local hypotheses due to the smaller effective sample size within each subgroup. In this paper, we fix $G_{\max}=1$ or $G_{\max}=2$, which achieve a reasonable balance and yield informative results. Optimally tuning $G_{\max}$ for a specific dataset is an interesting direction for future research.
As detailed in Appendix A2.2, this implementation can be extended to handle multiple binary covariates as well as non-binary ones. For non-binary covariates, we apply discretization, as the ability to detect meaningful subgroups and reject the corresponding null hypotheses generally depends on having a sufficient number of individuals in each subgroup. Additionally, Appendix A2.3 describes a specialized implementation with computational shortcuts designed for GWAS-like settings, characterized by a very large number $p$ of variables and a modest number $m$ of covariates.
@@ 3 Empirical Demonstrations
We demonstrate the performance of aLKF through two simulated settings and one real data analysis. The simulation settings are designed to capture the most practically relevant scenarios. The first, presented in Section 3.1, investigates variable selection across many environments ($m$) with a moderate number of variables ($p$). The second, described in Section 3.2, reflects GWAS-like conditions with high dimensionality ($p>n$) and fewer environments, serving as a stepping stone to the real GWAS data analysis in Section 3.3.
@@ 3.1 Simulation: Finding Highly Localized Important Variables
We generate synthetic data with $p=20$ binary variables and $m=80$ covariates, where the first 20 covariates are binary and the remaining 60 are continuous in $[0,1]$. Details on the joint distribution of these variables are in Appendix A4.
@@
Based on this simulation setup, valid knockoffs can be sampled independently conditional on the covariates from a Bernoulli distribution; see Appendix A4 for further details.
Given a sample of $n$ independent observations, our goal is to identify which variables are non-null within which subgroups, aiming for both power (many rejections) and precision (informative subgroups). We apply our method (aLKF) at the nominal FDR level $\alpha=0.1$, using the lasso-based approach for importance scores (Appendix A2.1) and learning data-driven partitions via a lasso model with interactions (Appendix A2.2), allowing at most two interactions per variable ($G_{\max}=2$). We compare aLKF against three benchmarks that are applicable within the knockoffs framework: (1) Global-KF, the knockoff filter of Candès et al. (2018), which controls the FDR but tests only population-wide hypotheses; (2) Split-LKF, a less powerful version of aLKF that splits the data, using half to learn the partition function and the other half to test local hypotheses, sacrificing power due to reduced sample size but maintaining FDR control; (3) Naive-LKF, an invalid version of aLKF that uses all data twice—to learn the partition and to test local hypotheses—without cloaking, failing to control the FDR due to selection bias.
@@ 3.2 Simulation: Searching for Gene-Environment Interactions
The following experiments simulate a search for gene-environment interactions in GWAS data, using real chromosome 21 genotypes from unrelated white non-British individuals in the UK Biobank. Working with real genetic variants introduces additional complexity due to strong collinearity among nearby variants on the same chromosome—a phenomenon known as linkage disequilibrium (LD). Because of LD, identifying the precise causal variant is often infeasible; instead, associations are typically detected at the level of broader genomic regions comprising multiple correlated SNPs. Following Sesia et al. (2021), we partition the genome into contiguous groups of SNPs in strong LD, which serve as the units of inference in place of individual variants. Intuitively, our method tests whether each group of SNPs is conditionally independent of the outcome within a given environment, conditional on all other SNPs and covariates. The formal definition of the local hypotheses for groups of SNPs is provided in Appendix A2.3.4.
The choice of SNP group size reflects a well-known trade-off between resolution and power: smaller groups enable more precise localization of associations but tend to reduce power due to strong correlations among nearby variants (Sesia et al., 2020, 2021; Gablenz and Sabatti, 2025). Sesia et al. (2021) introduce a hierarchy of LD-based groupings to support multi-resolution analyses, and Gablenz and Sabatti (2025) propose a method to aggregate results across resolutions while controlling the FDR. To maintain focus and clarity, we adopt a single resolution in this paper, with a median SNP group width of 3 kilobases (kb)—a scale that captures strong linkage disequilibrium while allowing variation in signal strength. Importantly, the group size does not affect the theoretical validity of our method. Further details on SNP grouping are provided in Appendices A2.3.4 and A5.
Because our hypotheses target groups of SNPs, we can use more powerful knockoffs that satisfy a weaker version of the exchangeability condition in Equation (7), which now only needs to hold under group-wise swapping. Using knockoffs designed for individual SNPs would reduce power unnecessarily. This issue has been addressed by Sesia et al. (2021), who developed an algorithm for generating valid knockoffs for SNP groups in GWAS data using hidden Markov models. We adopt their knockoffs in our analysis.
There are $14,733$ individuals in total and $p=8,832$ variants. We randomly sample 30% of the individuals ($n=4,420$) to imitate a more typical high-dimensional GWAS setting. We simulate two independent binary random covariates $Z_{1}$ and $Z_{2}$, independently across individuals from a Bernoulli($0.5$) distribution. The results of additional experiments using only one covariate can be found in Appendix A6. A continuous outcome $Y$ is simulated based on a linear model: $Y^{i}=\beta_{0}X^{i}+(\beta_{1}X^{i})*\mathbf{1}(Z_{1}^{i}=0)+\epsilon^{i}$, with $\beta_{0},\beta_{1}\in\mathbb{R}^{p}$, where $\epsilon^{i}\sim N(0,1)$ independently for each $i\in[n]$. Thus, $Z_{1}$ is used to create local effects, among the subgroup with $Z_{1}=0$, while $Z_{2}$ is unimportant. We randomly choose 1% of the genetic variants to be non-null, in the sense that they have a non-zero coefficient either globally (in $\beta_{0}$) or locally (in $\beta_{1}$). Additional results obtained using different proportions of non-null variants (0.05% and 2%) can be found in Appendix A6. Each non-null variant is randomly assigned as either a global or a local effect according to the desired proportions.
@@
Our goal is to identify which groups of SNPs are non-null within which subgroups of individuals, aiming for both high power and precise localization. We apply aLKF as described in Section 3.1, with modifications to account for SNP grouping as detailed in Appendix A2.3.4, and we constrain the data-driven partition function to select at most one interaction per variable ($G_{\max}=1$). In addition to the three benchmarks from Section 3.1Global-KF, Split-LKF, and Naive-LKF—we introduce a fourth, Fixed-LKF, a non-adaptive version of LKF described below. All methods in this comparison use the same SNP group definitions and the same group knockoffs from Sesia et al. (2021).